OpenStudy (anonymous):

What is 10^6 in standard form? A. 0.0000001 B. 0.000001 C. 1,000,000 D. 10,000,000

OpenStudy (alexandervonhumboldt2):

10^6=10*10*10*10*10*10=?

OpenStudy (anonymous):

\[10^6=10*10*10*10*10*10=?\]

OpenStudy (anonymous):

I think it is d but i am not sure

OpenStudy (anonymous):

So each time to multiply with 10, there comes an extra 0. So when you multiply 10 with 10 5 times, you add 5 zeroes. What would that be?

OpenStudy (anonymous):

so 10^5 would be 1000000

OpenStudy (anonymous):

No, \[10^5=10*10*10*10*10=100,000\]

OpenStudy (anonymous):

So \[10^6=100,000*10=?\]

OpenStudy (anonymous):

10,00000

OpenStudy (anonymous):

but since it is power of ten shouldn't there be 7 zeroes because the 10 already has one

OpenStudy (welshfella):

10^6 is 1 followed by 6 zeros

OpenStudy (anonymous):

so the answer is 1000000

OpenStudy (anonymous):

so then the answer will be c

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

my study guide says that this is powers of tens so i thought you have to include the zero from the ten thats why i added 7

OpenStudy (anonymous):

but thanks

OpenStudy (anonymous):

I have 1 more question do you think you can answer it